Assume that the production function for a country is given by and annual investment is given by the function where -0.3100, and that the yearly depreciation rate is 5.167%. Suppose that this year, the output in the country is l, and a neighbor countrys output is 50% higher. Calculate the time it would take for the countrys output to catch up with its neighbors output. Assume the neighbor countrys economy is neither growing nor shrinking. years
Answer #1

Answer

Y = K1/2 and I = yY = 0.31Y

=> I = 0.31*K1/2

Depreciation rate = 5.167% => Depreciation = 0.05167*K

dK/dt = Investment - depreciation = 0.31*K1/2 - 0.05167*K

=> Kt + 1 = Kt + 0.31*Kt1/2 - 0.05167*Kt ------------------------(1)

Initially Y = 1 => Y = K1/2 = 1

=> K = 1

We want Y = 1 + 50% of 1 = 1.5

=> K = 1.52 = 2.25

Hence we have to calculate time required to increase K from 1 to 1.5

Lets Use Excel to calculate t required to increase K from 1 to 2.25.

Initially K0 = 1

Using (1) , We have the following result :

C) 1.25833 2 1.541055548 3 1.846260908 4 2.172083877 5 2.516730034 6 2.878480951 7 3.255698525 8 3.646826766

We can see from the above table that time required to increase K from1 to 2.25 is approximately 5 years.

Hence, the correct answer is 5 Years